Owl shift summary, 18 Sep 1999

Valeri Jejer (jejer@fnal.gov)
Sat, 18 Sep 1999 07:32:17 -0500 (CDT)

Shift Summary for owl shift, 18/sep/99

On Shift: Theo, Valeri, Emmanuel

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 7.0
Hours of E799 data (special runs not included): 0.75
Hours of special runs (laser scans, muon runs...): 6.25

Accounting of lost time:

kfrends caused by ktevana errors (missing FFFF in Crate 53),
problems with resetting kumquats,
fixing STT and HC triggers
========================================================================
Runs ended this shift:
Run 14590-97 - - BEAM_799 no tapes written

Tapes brought to FCC this shift:
(including tapes for online split)

none

========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:

Cascade yield studies with 0", 10" and 20" Be absorber interrupted by
various hyperon trigger problems which were fixed.

========================================================================
Description of Shift:

Intermittent in the beginning of the shift beam (TeV injection problems)
got better by ~2 am and was stable around 7e12 after.

Theo and Emmanuel measured cascade yields with 0", 10" and 20" Be
absorber.
(see KTeV Log XXI pp.228-230)

run spill time comment
-------------------------------------------------------------------------
14591 00:24 0" abs. run for B10 trigger with HCCGE2 and no
prescale.
14592 00:51 initializing run w/10" abs. could not load
kumquats,
reloading them from ktrig window did not work.
Investigating the problem Emmanuel found that
STT cables were unplugged from the banana module.
Plugged them back. STT is working after that.
Kumquats reset, but kumq button was red at every
init not causing any DA problems.
Started this run with 10" abs., then switched to
no abs.
14593 03:00 Another run with no abs. after Emmanuel found that
HC cables were switched downstairs. Also,
there is a note in LogBook XX, p. 184 on this
problem that probably occured when people
swapped cables in a bad shaper/splitter crate.
By swapping the HC cables plugged into the
UCLA shapers and ADCs, the problem was solved.
Still we need to make an access to verify the
downstairs swap of the two HC cables.
14594 03:25 no abs. run for cascade yield studies.
14595 10" abs. run for cascade yield studies.
14596 20" abs. run for cascade yield studies.
14598 06:15 E799 run with 10" absorber and writing tapes.
========================================================================
Follow up Notes on Previous Shift

o Did not page Arun - were not down for ~>1 hour

========================================================================
Items to Follow up on Next Shift

o Page Arun if down for ~>1 hour

o Update the E799 Shift Check List on Primary Beam, etc.

-----------------------------------------------------------------------